OpenStudy (anonymous):

HELP! The measure of the angles of a triangle are represented by x+30, 4x+30, and 10x-30, the triangle must be: 1) isosceles 2) right 3) obtuse 4) scalene EXPLAIN PLEASE

OpenStudy (johnweldon1993):

Alright...so firstly you know that the angles of a triangle add up to 180 correct?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

So if you just had 3 angles you would add them...like 20 + 70 + 90 = 180 and you would be all set... same thing here....just add up all these functions and solve for x \[\large x + 30 + 4x + 30 + 10x - 30 = 180\] Now combine all your x terms...and combine all your constants and solve for x what do you get?

OpenStudy (anonymous):

x=10

OpenStudy (johnweldon1993):

Very good...now...to figure out your indivdual angles...you plug x = 10 into each of those equations... x + 30 becomes 10 + 30 = ? 4x + 30 becomes 4(10) + 30 = ? 10x - 30 becomes 10(10) - 30 = ?

OpenStudy (anonymous):

40, 70, 70

OpenStudy (johnweldon1993):

Perfect! which is a ____triangle?

OpenStudy (anonymous):

so it wouls be isosceles

OpenStudy (johnweldon1993):

Very good :)

OpenStudy (anonymous):

Thanks. I have a test on this AND proofs tomorrow.. :(

OpenStudy (johnweldon1993):

Yeah it is kind of confusing when you look at it but when you can step through it like we did here and think about what they ask it's not so bad :) Good luck on your test tomorrow!!

OpenStudy (anonymous):

Thank you, I doubt I will get any of the proofs right